OpenStudy (anonymous):

How can 2x2 = x2 + 4x + 8 be set up as a system of equations? y = x2 - 4 y = x2 + 2x + 8 y = x2 y = x2 + 4x + 8 y = 2x2 y = x2 + 4x + 8 y = x2 y = 2x + 8

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@ikram002p @paki

OpenStudy (anonymous):

@stezzynation

OpenStudy (anonymous):

well,look at C: y = 2x2 y = x2 + 4x + 8 so 2x2 = x2 + 4x + 8 as y =y